OpenStudy (anonymous):

Solve P = 4(3a - b) for a in terms of P and b. You must show all work

OpenStudy (anonymous):

P=4(3a-b)=12a-4b 12a=P+4b a=P/12+b/3

OpenStudy (anonymous):

but isn't it asking for b?

OpenStudy (anonymous):

It's asking you to solve for a in terms of P and b. If it was asking for b, it should have been written "Solve for b in terms of P and a."

OpenStudy (anonymous):

oh ok. Thank you so much!